Circular No. 6310

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 D. M. Cole, D. E. Vanden Berk, S. A. Severson, R. C. Nichol, M. C.

Miller, J. M. Quashnock, D. Q. Lamb, E. Bergeron, K. Gloria, and D. Long, University of Chicago, report: “We made near-infrared observations during Jan. 30.529–30.551 UT, and optical observations during Feb. 5.534–5.582, of a region of the sky that included the entire XTE error box (IAUC 6291, 6309) for GRO J1744–28, using the Astrophysical Research Consortium’s 3.5-m telescope at Apache Point Observatory. We have examined the position of the variable radio source reported on IAUC 6307 in stacked images in g (665 s), r (665 s), and K′ (510 s). In the stacked r image, we find an object at this position (within 1′′) that is 3.7 mag fainter in r than a star with photographic mag b_{j} = 18.26 in COSMOS/NRL scans of the U.K. Schmidt Southern Sky Survey. Assuming a color of b_{j}–r = +1.1, this implies an r = 21. Visual examination of an ESO copy of a U.K. Schmidt Survey red plate also reveals an object at this location at the limit of the plate. In the stacked g image, the faintest objects we confidently detect are g = 19.3 (using a color conversion of b_{j}–g = +0.7), but no object at all is visible at the position of the radio source. In the stacked K′ image, we find an object of mag 13 at this position. Examination of an earlier (1992) K-band map of the field kindly provided by M. Merrill and I. Gatley (NOAO) reveals an object at this location of comparable brightness. In these observations, the optical/NIR counterpart of the radio source does not appear significantly brighter than at earlier epochs. The r band results quoted on IAUC 6309, in conjunction with our own data, may indicate that the object has brightened in the past four days.”

Further to their item on IAUC 6309, T. Augusteijn et al. add: “A spectrum (range 400–1000 nm, resolution 2.8 nm) of the star coinciding with the radio source in the XTE error box of GRO J1744–28, taken with the New Technology Telescope on Feb. 9.332 UT, shows that this object is an M0 (± 1 subtype) star, likely a moderately reddened [E(B − V) ∼ 0.5 mag] star at a distance of 500–1000 pc. The spectrum shows no trace of Hα emission (upper limit on its equivalent width 0.05 nm). It is highly unlikely that this M star is the optical counterpart of GRO J1744–28. Therefore, if the radio source is connected with the x-ray source, the optical counterpart must be substantially fainter than V = 20.”

1996 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green